OpenStudy (anonymous):

The regular price of an item is $35. The item is on sale with a discount of $10. What is the discount rate? A. 0.286% B. 28.6% C. 35% D. 286%

OpenStudy (anonymous):

I got C.

OpenStudy (beccaboo333):

You're wrong.

OpenStudy (beccaboo333):

\[35\times.35\]

OpenStudy (beccaboo333):

That equals \[12.25\] not 10

OpenStudy (beccaboo333):

Okay, want to know how you'd get the correct answer?

OpenStudy (beccaboo333):

No.

OpenStudy (beccaboo333):

Do this for each one. \[35\times whatever options you have.\]

OpenStudy (anonymous):

Ur confusing me ;/

OpenStudy (beccaboo333):

@ganeshie8 will explain it.

OpenStudy (anonymous):

nvm listen to her i guess

OpenStudy (anonymous):

..ok..

ganeshie8 (ganeshie8):

^^thats a good idea to check the options quickly

ganeshie8 (ganeshie8):

or u could find the percentage : \(\large \frac{10}{35} \times 100\)

OpenStudy (anonymous):

ok

OpenStudy (beccaboo333):

Which ever option equals \[10\] or close to \[10\] (rounding) is your answer.

OpenStudy (beccaboo333):

Did you figure out your answer?

OpenStudy (anonymous):

yes

OpenStudy (beccaboo333):

Which option do you think it is?